The absorption spectra of ions have been used to identify the presence of the elements in the atmospheres of the sun and other stars. What is the energy of a photon (in J) that is absorbed by He+ ions, when an electron is excited from the Bohr orbit with n = 1 to the n = 4 state? The energy of an electron in the nth level is:
Answer #1

En   = -2.178*10^-18J*Z^2/n^2

Z = 2 for He^+

n = 1

E1 = -2.178*10^-18*(2)^2/(1)^2

E1    = -8.712*10^-18J

n = 4

E4 = -2.178*10^-18*(2)^2/(4)^2

E4   = -0.5445*10^-18J

E4-E1 = -0.5445*10^-18-( -8.712*10^-18)

             = 8.1675*10^-18J >>>>answer